National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1281–1300 of 2520Page 65 of 100
NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1281

National Defence Academy (NDA) 2020• Mathematics• Mathematics

What is the probability that February of a leap year selected at random, will have five Sundays?

  1. AOption (a)

    \(\dfrac{1}{5}\)

  2. BOption (b)

    \(\dfrac{1}{7}\)

  3. COption (c)

    \(\dfrac{2}{7}\)

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (b)B.

\(\dfrac{1}{7}\)

Why this answer is correct

Concept:

Total number of days in a leap year = 366

Total number of days in February = 29

Total no. of days in a week = 7

Total no. of weeks= 29/7= 4 weeks and 1 day

 

Calculation:

To find: Probability that a leap year selected at random will contain five Sundays

As we know, in a leap year in the february months there will be 4 weeks and 1 day. 

These 1 days can be:

{Sun}, {Mon}, {Tue}, {Wed}, {Thu}, {Fri}, {Sat} 

Total outcomes = 7

Favourable outcomes = {Sun} = 1

Hence the probability that a leap year selected at random will contain five Sundays = 1/7

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1282

National Defence Academy (NDA) 2020• Mathematics• Mathematics

The arithmetic mean of 100 observation is 40. Later, it was found that an observation '53' was wrongly read as '83'. What is the correct arithmetic mean?

  1. AOption (a)

    39.8

  2. BOption (b)

    39.7

  3. COption (c)

    39.6

  4. DOption (d)

    39.5

View answer & explanation

Correct answer

Answer: (b)B.

39.7

Why this answer is correct

Concept:

The mean is equal to the sum of all the values in the data set divided by the number of values in the data set.

Calculations:

Given, the arithmetic mean of 100 observation is 40.

The mean is equal to the sum of all the values in the data set divided by the number of values in the data set.

⇒ Sum of 100 observation = 100 \(\times ;40\) = 4000

Later, it was found that an observation '53' was wrongly read as '83

⇒ Sum of 100 observation = 4000 - 83 + 53 = 3970

New Mean = \(\dfrac {3970}{100}\)= 39.70

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1283

National Defence Academy (NDA) 2020• Mathematics• Mathematics

A husband and wife appear in an interview for two vacancies for the same post. The probability of the husband's selection is \(\dfrac{1}{7}\) and that of the wife's selection is \(\dfrac{1}{5}\). If the events are independent, then the probability of which one of the following is \(\dfrac{11}{35} \ ?\)

  1. AOption (a)

    At least one of them will be selected

  2. BOption (b)

    Only one of them will be selected

  3. COption (c)

    None of them will be selected

  4. DOption (d)

    Both of them will be selected

View answer & explanation

Correct answer

Answer: (a)A.

At least one of them will be selected

Why this answer is correct

Calculations:

Given , a husband and wife appear in an interview for two vacancies for the same post.

Let A= be the event that the husband is selected.

B = be the event that the wife is selected. 

P(A) = \(\dfrac{1}{7}\) and P(B) = \(\dfrac{1}{5}\)

Now, 

A̅ = be the event that the husband is not selected.

B̅  = be the event that the husband is not selected.

P(A̅) = 1 - P(A) = \(\dfrac{6}{7}\) and P(B̅) = 1 - P(B) = \(\dfrac{4}{5}\)

P(At least one of them will be selected) = P(A) P(B̅) + P(A̅) P(B) + P(A) P(B)

= \(\frac17 \times \frac 45 + \frac67 \times \frac 15 + \frac17 \times \frac 15 \)

= \(\dfrac{11}{35} \ \)

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1284

National Defence Academy (NDA) 2020• Mathematics• Mathematics

A dealer has a stock of 15 gold coins out of which 6 are counterfeits. A person randomly picks 4 of the 15 gold coins. What is the probability that all the coins picked will be counterfeits?

  1. AOption (a)

    \(\dfrac{1}{91}\)

  2. BOption (b)

    \(\dfrac{4}{91}\)

  3. COption (c)

    \(\dfrac{6}{91}\)

  4. DOption (d)

    \(\dfrac{15}{91}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\dfrac{1}{91}\)

Why this answer is correct

Concept:

Product rule:

If a probability of event A happening is x and probability of event B happening is y then the probability of happening event s A and B is xy.

Probability:

If S is a sample space and A is a favourable event then the probability of A is given by:

\(\rm P=\dfrac{n(A)}{n(S)}\)

Combination formula:

The number of possible selection (combinations) of r objects out of n objects is given by:

\(\begin{align*} \rm ^nC_r &=\rm \dfrac{n!}{(n-r)!r!} \end{align*}\)

 

Calculation:

First we will decide the sample space.

The sample space will contain all possible combinations where we select 4 coins out of 15 coins.

Therefore,

\(\begin{align*} ^{15}C_4 &= \dfrac{15!}{11!\times 4!}\ &= \dfrac{12\times13\times 14\times 15}{4\times 3\times 2\times 1}\ &= 91\times 15 \end{align*}\)

Now the favourable event is selecting 4 coins out of the counterfeit 6 coins which is given as follows:

\(\begin{align*} ^6C_4 &= \dfrac{6!}{2!\times 4!}\ &= \dfrac{5\times 6}{2\times 1}\ &= 15 \end{align*}\)

Therefore, the required probability is:

\(\begin{align*} P &= \dfrac{15}{91\times 15}\ &= \dfrac{1}{91} \end{align*}\)

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1285

National Defence Academy (NDA) 2020• Mathematics• Mathematics

A committee of 3 is to be formed from a group of 2 boys and 2 girls. What is the probability that the committee consists of 2 boys and 1 girl?

  1. AOption (a)

    \(\dfrac{2}{3}\)

  2. BOption (b)

    \(\dfrac{1}{4}\)

  3. COption (c)

    \(\dfrac{3}{4}\)

  4. DOption (d)

    \(\dfrac{1}{2}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\dfrac{1}{2}\)

Why this answer is correct

Concept:

Let A be any event in sample space S, then P(A) = \(\rm\dfrac{n(A)}{n(S)}\)

Calculations:

A committee of 3 is to be formed from a group of 2 boys and 2 girls.

Here are two boys B1 and B2 and two girls G1 and G2.

Sample space = S = {B1B2G1 , B1B2G2, G1G2B1, G1G2B2}

⇒n (S) = 4

A: The event that the committee consists of 2 boys and 1 girl

A = {B1B2G1 , B1B2G2}

⇒n(A) = 2

probability that the committee consists of 2 boys and 1 girl = P(A) = \(\rm\dfrac{n(A)}{n(S)}\)= \(\rm \dfrac 2 4 \)

Hence, A committee of 3 is to be formed from a group of 2 boys and 2 girls, the probability that the committee consists of 2 boys and 1 girl \(\dfrac 12 \)

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1286

National Defence Academy (NDA) 2020• Mathematics• Mathematics

In a lottery of 10 tickets numbered 1 to 10, two tickets are drawn simultaneously. What is the probability that both the tickets drawn have prime numbers?

  1. AOption (a)

    \(\dfrac{1}{15}\)

  2. BOption (b)

    \(\dfrac{1}{2}\)

  3. COption (c)

    \(\dfrac{2}{15}\)

  4. DOption (d)

    \(\dfrac{1}{5}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\dfrac{2}{15}\)

Why this answer is correct

Concept:

Let A and B are independent events, then 

P(A \(\rm\cap\) B)= P(A). P(B)

Calculations:

Give, total number of tickets = 10

Number of tickets drawn = 2

Prime number between 1 to 10 = {2, 3, 5, 7} = 4

A : The event of getting a prime number in the first draw

B : The event of getting a prime number in the second draw

The event A and B are independent events.

P(A) = Probability of getting a prime number in the first draw = \(\rm \dfrac{4}{10} = \dfrac {2}{5}\)

Number of remaining cards in pack = 9

P(B) = Probability of getting a prime number in the second draw = \(\rm \dfrac{3}{9} = \dfrac {1}{3}\)

The probability that both the tickets drawn have prime numbers P(A \(\rm\cap\) B)= P(A). P(B)

= \(\rm \dfrac{2}{5} \times \dfrac {1}{3}\)

= \(\rm\dfrac {2}{15}\)

Hence, In a lottery of 10 tickets numbered 1 to 10, two tickets are drawn simultaneously. then the probability that both the tickets drawn have prime numbers is \(\rm\dfrac {2}{15}\)

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1287

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Let X and Y represent prices (in Rs) of a commodity in Kolkata and Mumbai respectively. It is given X̅ = 65, Y̅ = 67, σX = 2.5, σY = 3.5 and r(X, Y) = 0.8. What is the equation of regression of Y on X ?

  1. AOption (a)

    Y = 0.175X - 5

  2. BOption (b)

    Y = 1.12X - 5.8

  3. COption (c)

    Y = 1.12X - 5

  4. DOption (d)

    Y = 0.17X + 5.8

View answer & explanation

Correct answer

Answer: (b)B.

Y = 1.12X - 5.8

Why this answer is correct

Concept:

Equation of regression of Y on X:

The equation of regression of Y on X with given values of \(\rm \bar X, \bar Y,\sigma_{x}\mbox{ and }\sigma_{y}\) is given as follows:

\(\rm \rm \dfrac{y-\bar Y}{\sigma_y} =r_{xy} \dfrac{x-\bar X}{\sigma_x} \)

 

Calculation:

It is given that X̅ = 65, Y̅ = 67, σX = 2.5, σY = 3.5 and r(X, Y) = 0.8 .

Using the equation of regression of Y on X we write:

\(\rm \rm \rm \frac{y-\bar Y}{\sigma_y} =r_{xy} \frac{x-\bar X}{\sigma_x} \ \left(\frac{y-67}{3.5}\right) =0.8 \times \left(\frac{x-65}{2.5}\right)\ y - 67 = 1.12(x-65)\y-67=1.12x-72.8\y=1.12x-5.8\)

Therefore, the required equation is y = 1.12x - 5.8

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1288

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Consider a random variable X which follows Binomial distribution with parameters n = 10 and \(\rm p = \dfrac{1}{5}\). Then Y = 10 - X follows Binomial distribution with parameters n' and p' respectively given by

  1. AOption (a)

    \(5, \dfrac{1}{5}\)

  2. BOption (b)

    \(5, \dfrac{2}{5}\)

  3. COption (c)

    \(10, \dfrac{3}{5}\)

  4. DOption (d)

    \(10, \dfrac{4}{5}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(10, \dfrac{4}{5}\)

Why this answer is correct

Concept:

Binomial distribution:

The binomial distribution for P(X = r) with parameters n and p is given as follows:

\(\rm P(X=r) = ^nC_rp^rq^{n-r}\)

Where q is the complementary event.

Also, the property of any random variable X,

E(aX + b)  = aE(X) + b

V(aX + b)  = (a2)V(X) 

Calculation:

Given: n = 10 and \(\rm p = \dfrac{1}{5}\)

Therefore, q = 1 - p 

\(\rm q = \dfrac{4}{5}\)

For random variable X, 

Mean = E(X) = np = 10 × (1/5) = 2

Variance = V(X) = npq = 10 × (1/5) ×(4/5)  = 8/5

Since, Y = 10 - X

⇒ E(Y) = E(10 - X)

⇒ E(Y) = 10 - E(X) = 10 - 2

⇒ 8 = n'p'      ....(1)

Also, V(Y) = V(10 - X) = 0 + (-1)2 V(X)

⇒ 8/5 = n'p'q'      ....(2)

Solving 1 and 2, we will get n', p', and q' which will be,

⇒ 8/5 = 8'q'   

⇒ q' = 1/5   

⇒ p' = 4/5   

Now,

⇒ 8 = n'  (4/5)

⇒ n' = 10

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1289

National Defence Academy (NDA) 2020• Mathematics• Mathematics

If A and B are two events such that P(A) = 0.6, P(B) = 0.5 and P(A ∩ B) = 0.4, then consider the following

Statements:

  1. P(A̅ ∪ B) = 0.9
  2. P(B̅ | A̅) = 0.6

Which of the statements is / are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (d)D.

Neither 1 nor 2

Why this answer is correct

Concept:

Probability of complementary event:

Let A be an event such that \(\rm \bar A\) is the complementary event then \(\rm P(A)=1-P(\bar A)\)  .

Probability of union of two events:

Let A and B be two events then the following relation holds:

\(\rm P(A\cup B)=P(A)+P(B)-P(A\cap B)\)

Conditional Probability :

For two events the conditional probability is given by:

\(\rm P(A|B)=\dfrac{P(A\cap B)}{P(B)}\)

 

Calculation:

It is given that \(\rm P(A)=0.6,P(B)=0.5\mbox{ and }P(A\cap B)=0.4\) .

We can calculate the probability of complementary events as follows \(\rm P(\bar A)=0.4,P(\bar B)=0.5\mbox{ and }P(\overline{A\cap B})=0.6\).

We know that 

\(\rm P(\bar A\cup B)=P(\bar A)+P(B)-P(\bar A\cap B)\)

Note that \(\rm P(\bar A\cap B)=P(B)-P(A\cap B)\) .

Therefore in the given case:

\(\rm P(\bar A\cap B) =0.5-0.4\ = 0.1 \)

Using this we write:

\(\rm P(\bar A\cup B) = 0.4+0.5-0.1\ = 0.8 \)

Therefore, the first statement is wrong.

Using the formula for conditional probability we write:

\(\rm P(\bar B|\bar A)=\dfrac{P(\bar A\cap \bar B)}{P(\bar A)}\)

We know that \(\rm P(\bar A\cap \bar B)=P(\overline{A\cup B})\) .

Using the given data we will first determine the value of \(\rm P(A\cup B)\).

\(\rm P(A\cup B) = P(A)+P(B)-P(A\cap B)\ = 0.6 + 0.5 - 0.4\ = 0.7 \)

Now using the complementary event formula we get:

\(\rm P(\overline{A\cup B}) = 1 - P(A\cup B)\ = 1 - 0.7\ = 0.3 \)

Substituting in the formula for the conditional formula:

\(\rm P(\bar B|\bar A) =\dfrac{P(\bar A\cap \bar B)}{P(\bar A)}\ = \dfrac{0.3}{0.4}\ = 0.75\)

Therefore, the second statement is also wrong.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1290

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Three cooks X, Y and Z bake a special kind of cake, and with respective probabilities 0.02, 0.03 and 0.05, it fails to rise. In the restaurant where they work, X bakes 50%, Y bakes 30% and Z bakes 20% of cakes. What is the proportion of failures caused by X?

  1. AOption (a)

    \(\dfrac{9}{29}\)

  2. BOption (b)

    \(\dfrac{10}{29}\)

  3. COption (c)

    \(\dfrac{19}{29}\)

  4. DOption (d)

    \(\dfrac{28}{29}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\dfrac{10}{29}\)

Why this answer is correct

Concept:

Consider, 

F : The event that the cake fails to raise.

A : The event that Cook X  bake a special kind of cake

B: The event that Cook Y  bake a special kind of cake

C : The event that Cook Z  bake a special kind of cake

The proportion of failures caused by X is given by P(A|F).

P(A|F) = \(\rm \dfrac {P(F|A)P(A)}{P(F|A)P(A)+P(F|B)P(B)+P(F|C)P(C)}\)

Calculations:

Consider, 

F : The event that the cake fails to raise.

A : The event that Cook X  bake a special kind of cake

B: The event that Cook Y  bake a special kind of cake

C : The event that Cook Z  bake a special kind of cake

Three cooks X, Y and Z bake a special kind of cake, and with respective probabilities 0.02, 0.03 and 0.05, it fails to rise.

P (F|A) = 0.02

P (F|B) = 0.03

P (F|C) = 0.05

In the restaurant where they work, X bakes 50%, Y bakes 30% and Z bakes 20% of cakes. 

P(A) = 0.5

P(B) = 0.3

P(C) = 0.2

To find : The proportion of failures caused by X is given by P(A|F).

P(A|F) = \(\rm \dfrac {P(F|A)P(A)}{P(F|A)P(A)+P(F|B)P(B)+P(F|C)P(C)}\)

Substitute the values, we get

P(A|F) = \(\rm \dfrac {(0.02)(0.5)}{(0.02)(0.5)+(0.03)(0.3)+(0.05)(0.2)}\)

P(A|F) = 34.48 = \(\dfrac{10}{29}\)

Hence, three cooks X, Y and Z bake a special kind of cake, and with respective probabilities 0.02, 0.03 and 0.05, it fails to rise. In the restaurant where they work, X bakes 50%, Y bakes 30% and Z bakes 20% of cakes. The proportion of failures caused by X is \(\dfrac{10}{29}\)

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1291

National Defence Academy (NDA) 2020• Mathematics• Mathematics

What is the length of the diameter of the sphere whose centre is at (1, -2, 3) and which touches the plane 6x - 3y + 2z - 4 = 0?

  1. AOption (a)

    1 unit

  2. BOption (b)

    2 units

  3. COption (c)

    3 units

  4. DOption (d)

    4 units

View answer & explanation

Correct answer

Answer: (d)D.

4 units

Why this answer is correct

Concept:

\(\begin{array}{l} \text { Distance of point } \rm \left (x_{1}, y_{1}, z_{1}\right) \text { from } a x+b y+c z+d=0 \text { is, D} =\rm \left|\frac{a x_{1}+b y_{1}+c z_{1}+d}{\sqrt{a^{2}+b^{2}+c^{2}}}\right| \end{array}\)

 

Calculation:

Here, centre of sphere is at (1, -2, 3) and equation of plane 6x - 3y + 2z - 4 = 0

So the distance from th point (1, -2, 3) to 6x - 3y + 2z - 4 = 0

\(\begin{array}{l} \rm r=\left|\frac{(1 \times 6)+(-2\times (-3))+(3 \times 2)-4}{\sqrt{6^{2}+(-3)^{2}+2^{2}}}\right| \ \rm r=\left|\frac{6+6+6-4}{\sqrt{49}}\right|=\frac{14}{7}\=2 \end{array}\)

Length of diameter = 2(2) = 4 units  

Hence, option (4) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1292

National Defence Academy (NDA) 2020• Mathematics• Mathematics

What is the perpendicular distance from the point (2, 3, 4) to the line \(\rm \frac{x-0}{1}=\frac{y-0}{0}=\frac{z-0}{0} \ ?\)

  1. AOption (a)

    6 units

  2. BOption (b)

    5 units

  3. COption (c)

    3 units

  4. DOption (d)

    2 units

View answer & explanation

Correct answer

Answer: (b)B.

5 units

Why this answer is correct

Concept:

Dot product of two perpendicular lines is zero.

Distance between two points (x1, y1, z1) and (x2, y2, z2) is given by, \(\rm \sqrt{(x_2-x_1)^2+(y_2-y_1)^2+(z_2-z_1)^2}\)

 

Calculation:

Let M be the foot of perpendicular drawn from the the point P(2, 3, 4)

Let , \(\rm \dfrac{x-0}{1}=\dfrac{y-0}{0}=\dfrac{z-0}{0} =k\)

x = k, y  = 0, z = 0

So M = (k, 0, 0)

Now direction ratios of PM = (2 - k, 3 - 0, 4 - 0) = (2- k, 3, 4) and direction ratios of given line are 1, 0, 0

PM is perepedicular to the given line so,

(2 - k) (1) + 3(0) + 4 (0) = 0

∴ k = 2

M = (2, 0, 0)

Perpendicular distance PM =

 \(\rm \sqrt {(2-2)^2+(0-3)^2+(0-4)^2}\ =\sqrt{9+16}\ =5\)

Hence, option (2) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1293

National Defence Academy (NDA) 2020• Mathematics• Mathematics

If a line has direction ratios < a + b, b + c, c + a >, then what is the sum of the squares of its direction cosines?

  1. AOption (a)

    (a + b + c)2

  2. BOption (b)

    2 (a + b + c)

  3. COption (c)

    3

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (d)D.

1

Why this answer is correct

Concept:

If l, m, n are direction cosines and a, b, c are direction ratios, then \(\rm l=\frac{a}{\sqrt{a^{2}+b^{2}+c^{2}}}, m=\frac{b}{\sqrt{a^{2}+b^{2}+c^{2}}} \text { and } n=\frac{c}{\sqrt{a^{2}+b^{2}+c^{2}}}\)

Calculation:

A line has direction ratios a + b, b + c, c + a

Let, direction cosines be l, m, and n 

So,

\(\rm l=\frac{a+b}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}}\)

\(m=\frac{b+c}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}} \)

\(n=\frac{c+a}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}}\)

\(\rm l^2+m^2+n^2=\rm [\frac{a+b}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}}]^2+[\frac{b+c}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}}]^2 +[\frac{c+a}{\sqrt{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}}]^2\)

\(\ =[\frac{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}{(a+b)^{2}+(b+c)^{2}+(c+a)^{2}}]\)

= 1

Hence, option (4) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1294

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Into how many compartments do the coordinate planes divide the space?

  1. AOption (a)

    2

  2. BOption (b)

    4

  3. COption (c)

    8

  4. DOption (d)

    16

View answer & explanation

Correct answer

Answer: (c)C.

8

Why this answer is correct

Explanation:

Two-dimensional coordinate axes divide the plane into four quadrants; 

Three-dimensional coordinate axes will divide the space into eight "compartments" known as octants.

Let (x, y, z) be an ordered triad. Take the point L of coordinate x on \(\overset \leftrightarrow{X'X}\) , the point M of coordinate y on \(\overset \leftrightarrow{Y'Y}\) and the point N of coordinate z on \(\overset \leftrightarrow{Z'Z}\). Through the point L, M, N draw planes π1 , π2 , π3  perpendicular to the coordinate axes \(\overset \leftrightarrow{X'X}\),\(\overset \leftrightarrow{Y'Y}\), \(\overset \leftrightarrow{Z'Z}\) respectively.

The three mutually perpendicular planes π1 , π2 , π3 intersect at a unique point of P.

Since P ϵ π1 , the projection of P on \(\overset \leftrightarrow{X'X}\) is L and hence the x - coordinate of P is equal to L. Similarly the y - coordinate of P is equal to the y - coordinate of M and the z -coordinate of p is equal to the z -coordinate of N. The coordinates of P are x, y, z in that order. Thus for the ordered triad (x, y, z) we have a unique point P in space.

Hence a one-to-one correspondence is established between the set of points in space and the set of ordered triads of real numbers. This space is called 3D space or R3 space.

The three coordinate planes divide the space into eight compartments, each of which is called an octant.

Octant with\(\overrightarrow{OX}\)\(\overrightarrow{OX'}\)\(\overrightarrow{OX'}\)\(\overrightarrow{OX}\)\(\overrightarrow{OX}\)\(\overrightarrow{OX'}\)\(\overrightarrow{OX'}\)\(\overrightarrow{OX}\)
founding\(\overrightarrow{OY}\)\(\overrightarrow{OY}\)\(\overrightarrow{OY'}\)\(\overrightarrow{OY'}\)\(\overrightarrow{OY}\)\(\overrightarrow{OY}\)\(\overrightarrow{OY'}\)\(\overrightarrow{OY'}\)
lines\(\overrightarrow{OZ}\)\(\overrightarrow{OZ}\)\(\overrightarrow{OZ}\)\(\overrightarrow{OZ}\)\(\overrightarrow{OZ'}\)\(\overrightarrow{OZ'}\)\(\overrightarrow{OZ'}\)\(\overrightarrow{OZ'}\)
NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1295

National Defence Academy (NDA) 2020• Mathematics• Mathematics

What is the equation of the plane which cuts an intercept 5 units on the z-axis and is parallel to xy-plane?

  1. AOption (a)

    x + y = 5

  2. BOption (b)

    z = 5

  3. COption (c)

    z = 0

  4. DOption (d)

    x + y + z = 5

View answer & explanation

Correct answer

Answer: (b)B.

z = 5

Why this answer is correct

Concept:

Plane parallel to xy- plane so, x-intercept = y-intercept = 0

 

Calculation:

Here, z-intercept = 5, and parallel to xy-plane

∴Equation of plane: z = 5

Hence, option (2) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1296

National Defence Academy (NDA) 2020• Mathematics• Mathematics

If â is a unit vector in the xy-plane making an angle 30° with the positive x-axis, then what is â equal to?

  1. AOption (a)

    \(\rm \dfrac{\sqrt{3}\hat{i}+\hat{j}}{2}\)

  2. BOption (b)

    \(\rm \dfrac{\sqrt{3}\hat{i}-\hat{j}}{2}\)

  3. COption (c)

    \(\rm \dfrac{\hat{i}+\sqrt{3}\hat{j}}{2}\)

  4. DOption (d)

    \(\rm \dfrac{\hat{i}-\sqrt{3}\hat{j}}{2}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\rm \dfrac{\sqrt{3}\hat{i}+\hat{j}}{2}\)

Why this answer is correct

Concept:

If vector making an angle θ with x-axis in xy plane, is given by, \(\rm cos \theta \hat i+sin\theta \hat j\)

 

Calculation:

Here, â is a unit vector in the xy-plane making an angle 30° with the positive x-axis

 \(\rm \vec a=cos 30 \hat i + sin 30 \hat j\ =\frac{\sqrt3}{2}\hat i+\frac{1}{2}\hat j\ =\frac{\sqrt 3\hat i+\hat j}{2}\)

Hence, option (1) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1297

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Let A be a point in space such that \(\rm |\overrightarrow{OA}|=12\), where O is the origin. If \(\rm\overrightarrow{OA}\) is inclined at angles 45° and 60° with x-axis and y-axis respectively, then what is \(\rm \overrightarrow{OA}\) equal to ?

  1. AOption (a)

    \(\rm 6\hat{i}+6\hat{j}\pm \sqrt{2} \ \hat{k}\)

  2. BOption (b)

    \(\rm 6\hat{i}+6\sqrt{2}\ \hat{j}\pm \ 6\hat{k}\)

  3. COption (c)

    \(\rm 6\sqrt{2} \hat{i}+6\hat{j}\pm 6\hat{k}\)

  4. DOption (d)

    \(\rm 3\sqrt{2} \hat{i}+3\hat{j}\pm 6\hat{k}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\rm 6\sqrt{2} \hat{i}+6\hat{j}\pm 6\hat{k}\)

Why this answer is correct

Concept: 

\(\rm cos^2α +cos^2β+cos^2γ =1\), where α, β, γ are the angles which inclined with X, Y, and Z axis respectively.

 

Calculation: 

Here, OA is inclined at angles 45° and 60° with x-axis and y-axis respectively.

 \(\rm cos^245+cos^260+cos^2\gamma=1\ \Rightarrow \frac{1}{2}+\frac{1}{4}+cos^2\gamma=1\ \Rightarrow cos^2\gamma =1-\frac{3}{4} =\frac{1}{4}\ \ \Rightarrow cos \gamma=\pm\frac{1}{2}\ So, \ \gamma =\pm60 \)

\(\rm\overrightarrow{O A}=|OA|(\cos 45^{\circ} \hat i+\cos 60^{\circ} \hat j\pm\cos 60^{\circ} \hat{k})\ =12(\frac{1}{\sqrt2}\hat i+\frac{1}{2}\hat j\pm\frac{1}{2}\hat k)\ \)

= \(\rm 6\sqrt{2} \hat{i}+6\hat{j}\pm 6\hat{k}\)

Hence, option (3) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1298

National Defence Academy (NDA) 2020• Mathematics• Mathematics

Two adjacent sides of a parallelogram are 2î - 4ĵ + 5k̂ and î - 2ĵ - 3k̂. What is the magnitude of dot product of vectors which represent its diagonals? 

  1. AOption (a)

    21

  2. BOption (b)

    25

  3. COption (c)

    31

  4. DOption (d)

    36

View answer & explanation

Correct answer

Answer: (c)C.

31

Why this answer is correct

Concept:

Triangle Law of Vector Addition: Triangle law of vector addition states that when two vectors are represented as two sides of the triangle with the order of magnitude and direction, then the third side of the triangle represents the magnitude and direction of the resultant vector.

 

⇒\(\rm\vec R=\vec a+\vec b\)

Parallelogram law of vector addition: 

If two vectors are acting simultaneously at a point, then it can be represented both in magnitude and direction by the adjacent sides drawn from a point. Therefore, the resultant vector is completely represented both in direction and magnitude by the diagonal of the parallelogram passing through the point

\(\rm\vec c=\vec a+\vec b\)

 

Calculation:

Here, vector a and b are adjacent sides of parallelogram

a = 2î - 4ĵ + 5k̂ and b = î - 2ĵ - 3k̂.

Let vector c and d are digonals of parallelogram 

Using parallelogram law, 

\(\rm\vec c=\vec a+\vec b\)

= (2î - 4ĵ + 5k̂) + (î - 2ĵ - 3k̂)

=.\(\rm 3\hat i-6\hat j+2\hat k\)

Now using triangle law,

\(\rm \vec b+\vec d = \vec a\ \Rightarrow \vec d=\vec a-\vec b\)

= (2î - 4ĵ + 5k̂) - (î - 2ĵ - 3k̂)

= \(\rm \hat i-2\hat j+8\hat k\)

\(\rm \vec c. \vec d=(\rm 3\hat i-6\hat j+2\hat k).(\rm \rm \hat i-2\hat j+8\hat k)\ =3+12+16\ =31\)

Hence, option (3) is correct.

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1299

National Defence Academy (NDA) 2020• Mathematics• Mathematics

If \(\rm |\vec{a}\times \vec{b}|^2 + |\vec{a}\cdot \vec{b}|^2=144\) and \(\rm |\vec{a}|=4\), then what is \(\rm |\vec{b}|\) equal to?

  1. AOption (a)

    3

  2. BOption (b)

    4

  3. COption (c)

    6

  4. DOption (d)

    8

View answer & explanation

Correct answer

Answer: (a)A.

3

Why this answer is correct

Concept:

Let \(\rm\vec {a}\) and \(\rm \vec {b}\) be the two vectors,

Dot product of two vectors  is given by:  \(\rm \vec{a}.\vec{b} = |\vec{a}|.|\vec{b}| cos θ\)

Cross product of two vectors  is given by: \(\rm \vec{a}× \vec{b} = |\vec{a}|.|\vec{b}| \sin θ; \hat{n}\)

 

Calculation:

Given:\(\rm |\vec{a}× \vec{b}|^2 + |\vec{a}\cdot \vec{b}|^2=144\) and \(\rm |\vec{a}|=4\)

To Find: \(\rm |\vec{b}|\)

\(\rm |\vec{a}× \vec{b}|^2 + |\vec{a}\cdot \vec{b}|^2 = |\vec{a}|^2|\vec{b}|^2 \sin^2 \theta +|\vec{a}|^2|\vec{b}|^2 \cos^2 \theta \=|\vec{a}|^2|\vec{b}|^2 [\sin^2 \theta+\cos^2 \theta]\=|\vec{a}|^2|\vec{b}|^2\)

Therefore, \(\rm |\vec{a}× \vec{b}|^2 + |\vec{a}\cdot \vec{b}|^2 = |\vec{a}|^2|\vec{b}|^2\)

⇒ 144 = 42 × \(\rm |\vec{b}|^2\)

⇒ \(\rm |\vec{b}|^2\) = 9

∴ \(\rm |\vec{b}|\) = 3

NDA Mathematics 6 Sep 2020 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1300

National Defence Academy (NDA) 2020• Mathematics• Mathematics

If the vectors \(\rm \vec{a}=2\hat{i}-3\hat{j}+\hat{k}, \vec{b}=\hat{i}+2\hat{j}-3\hat{k}\) and \(\rm \vec{c}=\hat{j}+p \hat{k}\) are coplanar, then what is the value of p?

  1. AOption (a)

    1

  2. BOption (b)

    -1

  3. COption (c)

    5

  4. DOption (d)

    -5

View answer & explanation

Correct answer

Answer: (b)B.

-1

Why this answer is correct

Concept:

 \(\begin{aligned} &\text { Let } \overrightarrow{\mathrm{a}}=\mathrm{a}_{1} \overrightarrow{\mathrm{i}}+\mathrm{b}_{1} \overrightarrow{\mathrm{j}}+\mathrm{c}_{1} \overrightarrow{\mathrm{k}}, \overrightarrow{\mathrm{b}}=\mathrm{a}_{2} \overrightarrow{\mathrm{i}}+\mathrm{b}_{2} \overrightarrow{\mathrm{j}}+\mathrm{c}_{2} \overrightarrow{\mathrm{k}} \text { and } \overrightarrow{\mathrm{c}}=\mathrm{a}_{3} \overrightarrow{\mathrm{i}}+\mathrm{b}_{3} \overrightarrow{\mathrm{j}}+\mathrm{c}_{3} \overrightarrow{\mathrm{k}} \text { be the three vectors }\ &\text { Condition for coplanarity: } \overrightarrow{\mathbf{a}} \cdot(\overrightarrow{\mathrm{b}} \times \overrightarrow{\mathrm{c}})=\left|\begin{array}{lll} \rm a_{1} & \mathrm{b}_{1} & \mathrm{c}_{1} \ \mathrm{a}_{2} & \mathrm{b}_{2} & \mathrm{c}_{2} \ \mathrm{a}_{3} & \mathrm{b}{3} & \mathrm{c}_{3} \end{array}\right|=0 \end{aligned}\)

Calculation:

Here, \(\rm \vec{a}=2\hat{i}-3\hat{j}+\hat{k}, \vec{b}=\hat{i}+2\hat{j}-3\hat{k}\) and \(\rm \vec{c}=\hat{j}+p \hat{k}\) are coplanar 

\(\begin{array}{l} \left|\begin{array}{ccc} 2 & -3 & 1 \ 1 & 2 & -3 \ 0 & 1 & p \end{array}\right|=0 \ \rm \Rightarrow 2(2 p+3)+3(p)+1(1)=0 \ \rm \Rightarrow 4p+6+3p+1=0 \ \rm \Rightarrow 7p+7=0 \ \rm \Rightarrow p=-1 \end{array}\)

Hence, option (2) is correct.

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.